# Analysis of Active/Inactive Patterns in the NHANES Data using Generalized Multilevel Functional Principal Component Analysis

### v2

@@ Abstract
Between 2011 and 2014 NHANES collected objectively measured physical activity data using wrist-worn accelerometers for tens of thousands of individuals for up to seven days. Here we analyze the minute-level indicators of being active, which can be viewed as binary (because there is an active indicator at every minute), multilevel (because there are multiple days of data for each study participant), functional (because within-day data can be viewed as a function of time) data. To extract within- and between-participant directions of variation in the data, we introduce Generalized Multilevel Functional Principal Component Analysis (GM-FPCA), an approach based on the dimension reduction of the linear predictor. Scores associated with specific patterns of activity are shown to be strongly associated with time to death. In particular, we confirm that increased activity is associated with time to death, a result that has been reported on other data sets. In addition, our method shows the previously unreported finding that maintaining a consistent day-to-day routine is strongly associated with a reduced risk of mortality (p-value $<0.001$) even after adjusting for traditional risk factors. Extensive simulation studies indicate that GM-FPCA provides accurate estimation of model parameters, is computationally stable, and is scalable in the number of study participants, visits, and observations within visits. R code for implementing the method is provided.
@@ 1 Introduction
Body-worn accelerometers allow high-resolution objective quantification of human physical activity. For example, both the US National Health and Nutrition Examination Survey (NHANES) and the UK Biobank study collected accelerometry data for tens of thousands of study participants. This paper focuses on the NHANES 2011-2014 accelerometry data, which was publicly released in December 2020. These data contain minute level summaries of physical activity for $13,603$ participants over seven days collected from wrist-worn accelerometers. Physical activity at each minute is summarized using a continuous measure called Monitor Independent Movement Summary units (MIMS) (John et al.,, 2019). In their original form, MIMS are difficult to interpret and translate into actionable information and physical activity recommendations. For this reason, it makes sense to first transform MIMS into measures that are easier to interpret. A few such transformations considered in the literature are: (1) binary data corresponding to active/inactive, walking/non-walking, or sleep/wake periods; (2) multinomial data corresponding to sedentary, low intensity physical activity (LIPA), and moderate to vigorous physical activity (MVPA); and (3) count data corresponding to the number of active seconds or steps in every minute. This paper focuses on the NHANES 2011-2014 binary active/inactive profile at the minute level over multiple days obtained by thresholding the continuous MIMS measures. However, the methods considered can be applied to any type of non-Gaussian data, including discrete data (e.g., binary, multinomial, counts) (Gaston et al.,, 2008; Kass and Ventura,, 2001; Kelly and Kass,, 2012; Sebastián-González et al.,, 2010; Sentürk et al.,, 2014; Swihart et al.,, 2015; Bothwell et al.,, 2022) or continuous data with strong departures from normality (Gaynanova et al.,, 2020; Staicu et al.,, 2018).
@@
To provide intuition about the data structure, Figure 1 displays the binary physical activity data for the $13,603$ subjects who participated in the NHANES 2011-2014 study. The binary active/inactive data is obtained by thresholding the MIMS data as $Z_{ij}(s)=\mathbf{1}_{\{W_{ij}(s)\geq 10.558\}}$, where $W_{ij}(s)$ corresponds to the $i^{\text{th}}$ individual’s MIMS measurement on day $j=1,\ldots,J_{i}$ at minute $s$. For most study participants, $J_{i}=7$, though some have a smaller number of days because some days were excluded for data quality purposes by NHANES. The threshold of $10.558$ for physical inactivity was suggested by Karas et al., (2022). The top panel shows the binary (black dots correspond to active minutes while white dots correspond to inactive minutes) activity profiles for all study participants. Data are sorted by age on the $y-$axis and are shown from midnight to midnight on the x-axis for Wednesday (top left) and Saturday (top right). Figure 1 reveals activity patterns of a representative sample of the US population. A quick inspection of the plots indicates that (1) people generally wake up later and go to bed later on Saturday than on Wednesday; (2) people in their $20$s tend to stay up past midnight and wake up later; and (3) children tend to be more inactive during the night and more active during the day than adults.
@@
To visualize the day-to-day activity patterns, the lower panels show data for two randomly selected study participants, a 17 year-old (left) and a 62 year-old (right). Each sub-panel corresponds to one day and displays both the binary data (grey dots) and smooth estimates of the probability of being active (blue curves) (Wood,, 2017). The plots indicate that there is substantial within- and between-day variability. Our goal is to model such non-Gaussian (e.g., binary) functional (e.g., minutes within a day) data observed at multiple time points (e.g., days).
@@
We are not aware of any existing functional data analysis method that can handle data of the size and structure illustrated in Figure 1. Indeed, the binary active/inactive profiles in the NHANES data exhibit the following characteristics: (1) a large number of study participants ($13,603$) and observations per day ($1,440$) (2) multiple days of observation (from $1$ to $7$); (3) substantial within- and between-day variability; (4) high within-participant correlation over the course of the day; and (5) non-Gaussian observations. To address the combination of these challenges we introduce generalized multilevel functional principal component analysis (GM-FPCA). Here “generalized” refers to the non-Gaussianity of the measurements, “multilevel” refers to the multiple days within each individual, and “functional” refers to the functional representation of data, where the basic unit of observation is a curve measured over some domain (in our example, time). In our data set, each functional observation is a 24-hour physical activity trajectory for a given participant on a specific day.
@@ 2 Literature review
Though we are not aware of any existing method that can handle non-Gaussian multilevel functional data of the size and complexity of the NHANES, there are several closely related ideas. GM-FPCA builds upon three methods in FDA: (1) functional principal components analysis (FPCA); (2) generalized functional principal component analysis (GFPCA) for non-Gaussian functional data; and (3) multilevel functional principal component analysis (MFPCA) for functional data observed at multiple visits. While there is a vast literature and reliable software for FPCA (Goldsmith et al.,, 2020; Jones and Rice,, 1992; Staniswalis and Lee,, 1998; Ramsay and Silverman,, 2005; Crainiceanu et al.,, 2023; Xiao et al., 2016a, ), GFPCA and MFPCA require a closer look to understand the current state of the literature.
Generalized functional principal components analysis (GFPCA) extends FPCA to non-Gaussian outcomes, such as binary or count data. More precisely, let $Z_{i}(s)$ denote the non-Gaussian functional observation for participant $i=1,\ldots,I$ at a point $s\in S$. The GFPCA model is
$g[E\{Z_{i}(s)\}]=\beta_{0}(s)+b_{i}(s)=\beta_{0}(s)+\sum_{l=1}^{\infty}\xi_{il}\phi_{l}(s)\;,$ (1)
@@
Several methods exist for estimating Model (1), starting with (Hall et al.,, 2008), who modeled $Z_{i}(s)$ using a latent Gaussian process with mean and covariance estimators constructed using the Taylor expansion of $g(\cdot)$. Bayesian methods for non-Gaussian functional data were proposed by (Gertheiss et al., 2017a, ; van der Linde,, 2009) for a variety of sampling scenarios. An Exponential Family Functional Principal Component Analysis (EFPCA) was proposed by Li et al., (2018), who used quadratic penalties on the scores and eigenfunctions of a low rank decomposition of the linear predictor. However, existing methods are either prohibitively slow for large datasets or are not accompanied by software. Notable exceptions are: (1) the variational Bayes approach for binary functional data (Wrobel et al.,, 2019) implemented in R via the registr::bfpca() function (Wrobel,, 2018; Wrobel and Bauer,, 2021); (2) the two-step GFPCA described in (Gertheiss et al., 2017b, ), which is implemented in the registr::gfpca_twoStep() function; and (3) a fast GFPCA method (Leroux et al.,, 2023) developed for any type of non-Gaussian data and deployed in the fastGFPCA package Wrobel, (2023). Briefly, fast GFPCA reduces the computational complexity by estimating the latent processes in small bins along the functional domain, recovering eigenfunctions $\{\phi_{l}(s)\}$ using the fast covariance estimation (FACE) algorithm (Xiao et al., 2016b, ), and estimating the scores $\{\xi_{il}\}$ using a generalized linear mixed-model framework conditional on the eigenfunctions $\{\phi_{l}(s)\}$.
@@
Multilevel functional principal components analysis (MFPCA) extends FPCA to the case when functional data are observed at multiple visits (days), but data are Gaussian. MFPCA was first proposed by Di et al., (2009), and is deployed the refund::mfpca.sc() function (Goldsmith et al.,, 2020). A much improved version of the software was deployed recently in the refund::mfpca.face() function based on the paper by Cui et al., (2022). Other developments include extensions to multivariate and multilevel functional data (Zhang et al.,, 2023).
@@
Extending these models to the case when both the outcomes are non-Gaussian and the functional data has a multilevel structure is not straightforward. However, several approaches have been proposed. In particular, Chen et al., (2013) introduced a reduced-rank method for the analysis of generalized functional data; however, they use a marginal approach, which does not estimate the random effects components that are central to our analysis. Scheipl et al., (2016) provides a framework for generalized function-on-scalar regression that can incorporate multilevel structure through subject-specific random effects; however, this approach is memory intensive and does not scale up well to the data application considered in this paper. For example, for a subset of $500$ study participants from NHANES, the model did not run using $64$GB of memory. With 128GB of memory, the method did run, but took 10 days to finish. To run the analysis on the entire data set may be possible by splitting larger data sets into smaller subsets and run analyses on each subset. However, exactly how these subset analyses would be combined across all study participants remains an open methods question that we cannot pursue in this paper. Goldsmith et al., (2015) proposed a generalized multilevel function-on-scalar regression based on FPCA of the underlying latent space. They used a Bayesian modeling approach based on posterior simulations. However, this approach is also not computationally feasible for the scale and complexity of the data considered here. As a point of reference, (Goldsmith et al.,, 2015) analyzed an accelerometer dataset of $583$ subjects with $5$ days per subject. After downsampling the data to $144$ points per day, the method still took $10$ days to run. This approach is not feasible for data sets of the size of NHANES, which contains tens of thousands of subjects, multiple days per subject, and $1440$ observations per day. Therefore, new computational approaches are necessary. This paper provides a new strategy for inference, describes its statistical and computational performance through extensive simulation studies, and applies it to the NHANES 2011-2014 data.
@@ 3 Methods
The observed data is of the form $\{s_{k},Z_{ij}(s_{k})\}$, where $s_{k}\in S$ ($k=1,...,K$) is a point in the domain $S$ and $Z_{ij}(s_{k})$ is the non-Gaussian observation for subject $i$ ($i=1,...,I$), visit $j$ ($j=1,...,J$), and at the point $s_{k}$. For example, in the dichotomized NHANES 2011-2014 data, $Z_{11}(0)$ is a binary variable indicating whether subject 1 was active at midnight on day 1. While data is only observed on a discrete grid $\{s_{k},\,k=1,...,K\}$, we assume that they are generated by a latent continuous process. Specifically, we assume that
$g[\mathop{\mathbb{E}}\{Z_{ij}(s)\}]=\beta_{0}(s)+a_{i}(s)+b_{ij}(s)\;,$ (2)
@@
$g[\mathop{\mathbb{E}}\{Z_{ij}(s)\}]=\beta_{0}(s)+\sum_{l=1}^{\infty}\xi_{il}\phi_{l}^{(1)}(s)+\sum_{m=1}^{\infty}\zeta_{ijm}\phi_{m}^{(2)}(s)\;.$ (3)
Fitting this model requires a “divide and conquer” approach, which partitions the problem into a sequence of smaller problems that are each individually tractable. First, we divide the data domain into $K$ overlapping bins centered at the grid points $s_{k}\in S$, $k=1,\ldots,K$. Second, we fit GLMMs in each data bin with both subject- and subject-visit random effects, and extract subject-visit estimates on the linear predictor scale. This step is crucial because it transforms non-Gaussian data into continuous predictors, upon which MFPCA developed for Gaussian data can be applied. Third, we conduct MFPCA on the estimates from the second step to obtain subject- and subject-visit level eigenfunction estimates. Fourth, a joint GLMM is fit conditional on the eigenfunctions estimated in the previous step. We now expound on each step and provide the necessary details.
@@
$g[E\{Z_{ij}(s\mid s\in S_{k})\}]=\beta_{0}^{*}(s_{k})+a_{i}^{*}(s_{k})+b_{ij}^{*}(s_{k})$ (4)
for each of the $K$ bins, $S_{1},...,S_{K}$. The notation $Z_{ij}(s\mid s\in S_{k})$ emphasizes that GLMMs are fit separately for each bin. The terms $\beta_{0}^{*}(s_{k})$, $a_{i}^{*}(s_{k})$, and $b_{ij}^{*}(s_{k})$ are the fixed effect intercept, subject random intercept, and subject-visit random intercept, respectively, for the bin $S_{k}$. The superscript $*{}^{*}$ distinguishes these “local” quantities from their global counterparts defined in Equation 2. This step only requires fitting a generalized linear mixed effects model with a fixed intercept and nested two-level random intercepts, and is agnostic to the software or the fitting method (e.g., Bayesian, frequentist). The approach naturally deals with missing data at the subject or visit level and borrows strengths across individuals with sparser sampling points or less information (high likelihood of zeros or one.) The estimated linear predictors $\widehat{\eta_{ij}^{*}}(s_{k})=\widehat{\beta}_{0}^{*}(s_{k})+\widehat{a_{i}^{*}}(s_{k})+\widehat{b_{ij}^{*}}(s_{k})$ are then used in Step 3.
Step 3: Decompose the $\widehat{\eta_{ij}^{*}}(s_{k})$ estimates for each bin using MFPCA (Di et al.,, 2009; Cui et al.,, 2022). We use the fast multilevel FPCA method (Cui et al.,, 2022) because it works seamlessly with high dimensional data. MFPCA produces smooth estimates of eigenfunctions at subject- and subject-visit levels, denoted by $\widehat{\phi}_{l}^{(1)}(s)$ ($l=1,...,L$) and $\widehat{\phi}_{m}^{(2)}(s)$ ($m=1,...,M$), respectively. The choice of $L$ and $M$ will be discussed in Section 3.2. These eigenfunctions will be used in Step 4.
Step 4: Given the estimated eigenfunctions $\{\widehat{\phi}_{l}^{(1)}(s)\}$ and $\{\widehat{\phi}_{m}^{(2)}(s)\}$, we need to estimate the scores for each subject-visit pair. Conceptually, this can be achieved by fitting a global GLMM that includes a random slope for each eigenfunction. This model uses observations from the entire domain $S$, hence the word “global” to differentiate it from the local models in Step 2. The global model is
$g[E\{Z_{ij}(s\mid\{\widehat{\phi}_{l}^{(1)}(s)\}_{l=1}^{L},\{\widehat{\phi}_{m}^{(2)}(s)\}_{m=1}^{M})\}]=\beta_{0}(s)+\sum_{l=1}^{L}\xi_{il}\widehat{\phi}_{l}^{(1)}(s)+\sum_{m=1}^{M}\zeta_{ijm}\widehat{\phi}_{m}^{(2)}(s).$ (5)
@@
$[Z_{ij}(s)\mid\zeta_{ijm},{\rm others}][\zeta_{ijm}\mid{\rm others}]\;,$

### v3

@@ Abstract
Between 2011 and 2014 NHANES collected objectively measured physical activity data using wrist-worn accelerometers for tens of thousands of individuals for up to seven days. Here we analyze the minute-level indicators of being active, which can be viewed as binary (because there is an active indicator at every minute), multilevel (because there are multiple days of data for each study participant), functional (because within-day data can be viewed as a function of time) data. To extract within- and between-participant directions of variation in the data, we introduce Generalized Multilevel Functional Principal Component Analysis (GM-FPCA), an approach based on the dimension reduction of the linear predictor. Scores associated with specific patterns of activity are shown to be strongly associated with time to death. Extensive simulation studies indicate that GM-FPCA provides accurate estimation of model parameters, is computationally stable, and is scalable in the number of study participants, visits, and observations within visits. R code for implementing the method is provided.
@@ 1 Introduction
Body-worn accelerometers allow high-resolution objective quantification of human physical activity. For example, both the US National Health and Nutrition Examination Survey (NHANES) and the UK Biobank study collected accelerometry data for tens of thousands of study participants. This paper focuses on the NHANES 2011-2014 accelerometry data, which was publicly released in December 2020. These data contain minute level summaries of physical activity for $13,603$ participants over seven days collected from wrist-worn accelerometers. Physical activity at each minute is summarized using a continuous measure called Monitor Independent Movement Summary units (MIMS) (John et al.,, 2019). In their original form, MIMS are difficult to interpret and translate into actionable information and physical activity recommendations. For this reason, it makes sense to first transform MIMS into measures that are easier to interpret. A few such transformations considered in the literature are: (1) binary data corresponding to active/inactive, walking/non-walking, or sleep/wake periods; (2) multinomial data corresponding to sedentary, low intensity physical activity (LIPA), and moderate to vigorous physical activity (MVPA); and (3) count data corresponding to the number of active seconds or steps in every minute. This paper focuses on the NHANES 2011-2014 binary active/inactive profile at the minute level over multiple days obtained by thresholding the continuous MIMS measures. However, the methods considered can be applied to any type of non-Gaussian data, including discrete data (e.g., binary, multinomial, counts) (Gaston et al.,, 2008; Kass and Ventura,, 2001; Kelly and Kass,, 2012; Sebastián-González et al.,, 2010; Sentürk et al.,, 2014; Swihart et al.,, 2015; Bothwell et al.,, 2022; Cui et al., 2022a, ) or continuous data with strong departures from normality (Gaynanova et al.,, 2020; Staicu et al.,, 2018).
@@
To provide intuition about the data structure, Figure 1 displays the binary physical activity data for the $13,603$ subjects who participated in the NHANES 2011-2014 study. The binary active/inactive data is obtained by thresholding the MIMS data as $Z_{ijk}=\mathbf{1}_{\{W_{ijk}\geq 10.558\}}$, where $W_{ijk}$ corresponds to the $i^{\text{th}}$ individual’s MIMS measurement on day $j=1,\ldots,J_{i}$ at minute $k$. For most study participants, $J_{i}=7$, though some have a smaller number of days because some days were excluded for data quality purposes by NHANES. The threshold of $10.558$ for physical inactivity was suggested by Karas et al., (2022). The top panel shows the binary (black dots correspond to active minutes while white dots correspond to inactive minutes) activity profiles for all study participants. Data are sorted by age on the $y-$axis and are shown from midnight to midnight on the x-axis for Wednesday (top left) and Saturday (top right). Figure 1 reveals activity patterns of a representative sample of the US population. A quick inspection of the plots indicates that (1) people generally wake up later and go to bed later on Saturday than on Wednesday; (2) people in their $20$s tend to stay up past midnight and wake up later; and (3) children tend to be more inactive during the night and more active during the day than adults.
@@
To visualize the day-to-day activity patterns, the lower panels show data for two randomly selected study participants, a 17 year-old (left) and a 62 year-old (right). Each sub-panel corresponds to one day and displays both the binary data (grey dots) and smooth estimates of the probability of being active (blue curves) (Wood,, 2017). The plots indicate that there is substantial within- and between-day variability. Our goal is to model such generalized (e.g., binary) functional (e.g., minutes within a day) data observed at multiple time points (e.g., days).
@@
We are not aware of any existing functional data analysis method that can handle data of the size and structure illustrated in Figure 1. Indeed, the binary active/inactive profiles in the NHANES data exhibit the following characteristics: (1) a large number of study participants ($13,603$) and observations per day ($1,440$) (2) multiple days of observation (from $1$ to $7$); (3) substantial within- and between-day variability; (4) high within-participant correlation over the course of the day; and (5) binary observations. To address the combination of these challenges we introduce generalized multilevel functional principal component analysis (GM-FPCA). Here “generalized” refers to the non-Gaussianity of the measurements, “multilevel” refers to the multiple days within each individual, and “functional” refers to the functional representation of data, where the basic unit of observation is a curve measured over some domain (in our example, time). In our data set, each functional observation is a 24-hour physical activity trajectory for a given participant on a specific day.
@@ 2 Literature review
Though we are not aware of any existing method that can handle generalized multilevel functional data of the size and complexity of the NHANES, there are several closely related ideas. GM-FPCA builds upon three methods in FDA: (1) functional principal components analysis (FPCA); (2) generalized functional principal component analysis (GFPCA) for generalized single-level functional data; and (3) multilevel functional principal component analysis (MFPCA) for functional data observed at multiple visits. While there is a vast literature and reliable software for FPCA (Goldsmith et al.,, 2020; Jones and Rice,, 1992; Staniswalis and Lee,, 1998; Ramsay and Silverman,, 2005; Crainiceanu et al.,, 2023; Xiao et al., 2016a, ), GFPCA and MFPCA require a closer look to understand the current state of the literature.
Generalized functional principal components analysis (GFPCA) extends FPCA to exponential-family outcomes such as binary or count data. More precisely, let $Z_{ik}$ denote the non-Gaussian functional observation for participant $i=1,\ldots,I$ at a point $s_{k}\in S$. The GFPCA model is
$g\Bigl{(}\mathop{\mathbb{E}}\bigl{[}Z_{ik}\mid\{\xi_{il}\}_{l=1}^{\infty}\bigr{]}\Bigr{)}=\beta_{0}(s_{k})+b_{i}(s_{k})=\beta_{0}(s_{k})+\sum_{l=1}^{\infty}\xi_{il}\phi_{l}(s_{k})\;,$ (1)
@@
Several methods exist for estimating Model (1), starting with Hall et al., (2008), who modeled $Z_{ik}$ using a latent Gaussian process with mean and covariance estimators constructed using the Taylor expansion of $g(\cdot)$. Bayesian methods for non-Gaussian functional data were proposed by Gertheiss et al., 2017a and van der Linde, (2009) for a variety of sampling scenarios. An Exponential Family Functional Principal Component Analysis (EFPCA) was proposed by Li et al., (2018), who used quadratic penalties on the scores and eigenfunctions of a low rank decomposition of the linear predictor. However, existing methods are either prohibitively slow for large datasets or are not accompanied by software. Notable exceptions are: (1) the variational Bayes approach for binary functional data (Wrobel et al.,, 2019) implemented in R via the registr::bfpca() function (Wrobel,, 2018; Wrobel and Bauer,, 2021); (2) the two-step GFPCA described in (Gertheiss et al., 2017b, ), which is implemented in the registr::gfpca_twoStep() function; and (3) a fast GFPCA method (Leroux et al.,, 2023) developed for any type of non-Gaussian data and deployed in the fastGFPCA package (Wrobel,, 2023). Briefly, fast GFPCA reduces the computational complexity by estimating the latent processes in small bins along the functional domain, recovering eigenfunctions $\{\phi_{l}(s)\}$ using the fast covariance estimation (FACE) algorithm (Xiao et al., 2016b, ), and estimating the scores $\{\xi_{il}\}$ using a generalized linear mixed-model framework conditional on the eigenfunctions $\{\phi_{l}(s)\}$.
@@
Multilevel functional principal components analysis (MFPCA) extends FPCA to the case when functional data are observed at multiple visits (days), but data are Gaussian. MFPCA was first proposed by Di et al., (2009), and is deployed in the refund::mfpca.sc() function (Goldsmith et al.,, 2020). A much improved version of the software was deployed recently in the refund::mfpca.face() function based on the paper by Cui et al., 2022b . Other developments include extensions to multivariate and multilevel functional data (Zhang et al.,, 2023).
@@
Extending these models to the case when both the outcomes are non-Gaussian and the functional data has a multilevel structure is not straightforward. However, several approaches have been proposed. In particular, Chen et al., (2013) introduced a reduced-rank method for the analysis of generalized functional data; however, they use a marginal approach, which does not estimate the random effects components that are central to our analysis. Scheipl et al., (2016) provides a framework for generalized function-on-scalar regression that can incorporate multilevel structure through subject-specific random effects; however, this approach is memory intensive and does not scale up well to the data application considered in this paper. For example, for a subset of $500$ study participants from NHANES, the model did not run using $64$GB of memory. With 128GB of memory, the method did run, but took 10 days to finish. To run the analysis on the entire data set may be possible by splitting larger data sets into smaller subsets and run analyses on each subset. However, exactly how these subset analyses would be combined across all study participants remains an open methods question that we cannot pursue in this paper. Goldsmith et al., (2015) proposed a generalized multilevel function-on-scalar regression based on FPCA of the underlying latent space. They used a Bayesian modeling approach based on posterior simulations. However, this approach is also not computationally feasible for the scale and complexity of the data considered here. As a point of reference, Goldsmith et al., (2015) analyzed an accelerometer dataset of $583$ subjects with $5$ days per subject. After downsampling the data to $144$ points per day, the method still took $10$ days to run. This approach is not feasible for data sets of the size of NHANES, which contains tens of thousands of subjects, multiple days per subject, and $1440$ observations per day. Therefore, new computational approaches are necessary. This paper provides a new strategy for inference, describes its statistical and computational performance through extensive simulation studies, and applies it to the NHANES 2011-2014 data.
@@ 3 Methods
The observed data is of the form $\{s_{k},Z_{ijk}\}$, where $s_{k}\in S$ ($k=1,...,K$) is a point in the domain $S$ and $Z_{ijk}$ is the non-Gaussian observation for subject $i$ ($i=1,...,I$), visit $j$ ($j=1,...,J$), and at the point $s_{k}$. For example, in the dichotomized NHANES 2011-2014 data, $Z_{111}$ is a binary variable indicating whether subject 1 was active at midnight on day 1. While data is only observed on a discrete grid $\{s_{k},\,k=1,...,K\}$, we assume that they are generated by a latent continuous process. Specifically, we assume that
$g[\mathop{\mathbb{E}}\{Z_{ijk}\}]=\beta_{0}(s_{k})+a_{i}(s_{k})+b_{ij}(s_{k})\;,$ (2)
@@
$g\Bigl{(}\mathop{\mathbb{E}}\bigl{[}Z_{ijk}\mid\{\xi_{il}\}_{l=1}^{\infty},\{\zeta_{ijm}\}_{m=1}^{\infty}\bigr{]}\Bigr{)}=\beta_{0}(s_{k})+\sum_{l=1}^{\infty}\xi_{il}\phi_{l}^{(1)}(s_{k})+\sum_{m=1}^{\infty}\zeta_{ijm}\phi_{m}^{(2)}(s_{k})\;.$ (3)
Fitting this model requires a “divide and conquer” approach, which partitions the problem into a sequence of smaller problems that are each tractable. First, we divide the data domain into $K$ overlapping bins centered at the grid points $s_{k}\in S$, $k=1,\ldots,K$. Second, we fit GLMMs in each data bin with both subject- and subject-visit random effects, and extract subject-visit estimates on the linear predictor scale. This step is crucial because it transforms non-Gaussian data into continuous predictors, upon which MFPCA developed for Gaussian data can be applied. Third, we conduct MFPCA on the estimates from the second step to obtain subject- and subject-visit level eigenfunction estimates. Fourth, a joint GLMM is fit conditional on the eigenfunctions estimated in the previous step. We now expound on each step and provide the necessary details.
@@
$g[\mathop{\mathbb{E}}\{Z_{ijk}\mid k:s_{k}\in S_{k}\}]=\beta_{0k}^{*}+a_{ik}^{*}+b_{ijk}^{*}$ (4)
for each of the $K$ bins, $S_{1},...,S_{K}$. The notation $\{Z_{ijk}\mid k:s_{k}\in S_{k}\}$ emphasizes that GLMMs are fit separately for each bin. The terms $\beta_{0k}^{*}$, $a_{ik}^{*}$, and $b_{ijk}^{*}$ are the fixed effect intercept, subject random intercept, and subject-visit random intercept, respectively, for the bin $S_{k}$. The superscript distinguishes these “local” quantities from their global counterparts defined in Equation 2. This step only requires fitting a generalized linear mixed effects model with a fixed intercept and nested two-level random intercepts, and is agnostic to the software or the fitting method (e.g., Bayesian, frequentist). The approach naturally deals with missing data at the subject or visit level and borrows strengths across individuals with sparser sampling points or less information (high likelihood of zeros or one). The estimated linear predictors $\widehat{\eta}_{ijk}^{*}=\widehat{\beta}_{0k}^{*}+\widehat{a}_{ik}^{*}+\widehat{b}_{ijk}^{*}$ are then used in Step 3.
Step 3: Decompose the $\widehat{\eta}_{ijk}^{*}$ estimates for each bin using MFPCA (Di et al.,, 2009; Cui et al., 2022b, ). We use the fast multilevel FPCA method (Cui et al., 2022b, ) because it works seamlessly with high dimensional data. MFPCA produces smooth estimates of eigenfunctions at subject- and subject-visit levels, denoted by $\widehat{\phi}_{l}^{(1)}(s)$ ($l=1,...,L$) and $\widehat{\phi}_{m}^{(2)}(s)$ ($m=1,...,M$), respectively. The choice of $L$ and $M$ will be discussed in supplementary materials Section S.2. These eigenfunctions will be used in Step 4.
Step 4: Given the estimated eigenfunctions $\{\widehat{\phi}_{l}^{(1)}(s)\}_{l=1}^{L}$ and $\{\widehat{\phi}_{m}^{(2)}(s)\}_{m=1}^{M}$, we need to estimate the population mean function and scores $\{\xi_{il}\}_{l=1}^{L}$ and $\{\zeta_{ijm}\}_{m=1}^{M}$. Conceptually, this can be achieved by fitting a global GLMM that includes a random slope for each eigenfunction. This model uses observations from the entire domain $S$, hence the word “global” to differentiate it from the local models in Step 2. The global model is
$g[\mathop{\mathbb{E}}\{Z_{ijk}\mid k:s_{k}\in S\}]=\beta_{0}(s_{k})+\sum_{l=1}^{L}\xi_{il}\widehat{\phi}_{l}^{(1)}(s_{k})+\sum_{m=1}^{M}\zeta_{ijm}\widehat{\phi}_{m}^{(2)}(s_{k}).$ (5)
@@
$[Z_{ijk}\mid\zeta_{ijm},{\rm others}][\zeta_{ijm}\mid{\rm others}]\;,$